# Voting-Based Semi-Parallel Proof-of-Work Protocol

### v1

@@ IV-B Incentive Attacks
In Appendix D, we explain the MDP model with the appropriate state and action space as well as the objective function for our protocol. In Figure 5, we present the results of the MDP models together with the results of Section II where we analyzed the relative rewards of the existing parallel PoW protocols. We pick $\gamma=0.5$ for the MDP models and truncate the results of Section II before they rise above the honest behavior for the sake of neatness in the figures. For Nakamoto consensus, we run the MDP model of [5] and the modified model described in in Appendix D is run for our protocol with $10^{-5}$ precision where $\alpha$ values range from $0.02$ to $0.48$ with a $0.02$ spacing between them.
@@
In our protocol, matching always gives a half reward when $(a,h,fork)=(1,1,relevant)$ (see Appendix D) whereas the adversarial reward in Nakamoto consensus increases with increasing $\gamma$ since matching succeeds more often. Thus, in comparison to the Nakamoto consensus, our protocol is more resilient against selfish mining attacks when $\gamma\geq 0.3$ and resilience increases even more as $\gamma$ increases further, which we plot in Appendix D. On the other hand, when $\alpha>0.29$ as $\gamma$ drops below $0.3$, Nakamoto consensus becomes slightly more resilient. For the relative ratio in other values of $\gamma$, we refer the reader to Appendix D.
@@ IV-C Transaction Fee Split Rule
In Appendix E, we consider such an attacker that acts against the rule of picking the ledger offering the most fees to increase its expected transaction fee revenue. We show that, as long as $r$ is small enough, such an attack is usually not profitable. Moreover, even if profitable, the effect on the transaction fee rewards received by honest miners is minimal.
V Discussion and Limitations
@@ Appendix B Why Relative Reward Metric?
To understand why relative reward metric is the correct choice and why the normalized reward metric of [28, 29] is an illusion, consider the adversarial (honest resp.) miners who hold $\alpha$ ($1-\alpha$ resp.) proportion of the hash power in the system. Let the total adversarial (honest resp.) reward per chain progress be denoted as $\mu$ ($\nu$ resp.). Clearly, an incentive compatible system that is not under an attack, should satisfy
@@
We will analyze this system in more detail. However, to give the gist of our argument in a few words, if such a system that is under attack is safe as claimed in [28, 29], then honest mining should be sustainable even if the adversary launches the attack. However, if honest mining is sustainable under the attack given the mining rewards and costs, than the adversarial mining is not only sustainable but also more profitable since adversary gets more relative rewards compared to its costs.
@@
$\frac{\nu_{attack}}{1-\rho}=\frac{\mu_{attack}}{\rho}<\frac{\mu_{attack}}{\alpha}.$ (15)
@@
$\frac{\mu_{attack}}{\rho}=\frac{\nu_{attack}}{1-\rho}>\frac{\nu_{attack}}{1-\alpha},$ (17)
@@ Appendix C Proof Withholding Attack for Tailstorm and DAG-style Voting
Figure 7: Tailstorm witholding attack
@@ C-B Simulations
Appendix D MDP Models for the Proposed Protocol
@@
To consider if this last case of expanded state space helps the adversary or not, we can first compare the relative rewards of MDP model of [5] and the reward split MDP we explained earlier. If the reward split MDP results in higher relative rewards for an $(\alpha,\gamma)$ pair, it makes no sense for the adversary to withhold the ledger with the most fees. On the other hand, if the reward split MDP results in less relative rewards, then it is in the interest of the adversary to launch this type of attack for initiator heights if it has a ledger offering the most fees. However, if it fails to override the honest branch eventually, it loses an extra coinbase rewards of the last proof height due to not sharing the ledger in time.
We ignore the reward loss and assume the adversary can always create a ledger with most fees offered, which is more than what the adversary could achieve. We implement this complicated situation where we have a state space of the form $(a,h,fork,p,l)$ with $l$ keeping track of the last common proof height. Even in the scenario where there is no reward loss, the the result is approximately a weighted average between the reward ratio of the work in [5] (with approximate weight $1/L$) and reward split MDP (with approximate weight (L1)/L(L-1)/L). This is expected since the last case arises in approximately $\frac{1}{L}$ of the proof heights, i.e., the initiator heights. On the other hand, for $\gamma<0.3$, the result is the same as the reward ratio of reward split MDP since it is more profitable to have parallel proofs in this case. As the gain is minor even when the adversarial losses are ignored, we omit the simulation with expanded state space as the computational burden becomes very large with increasing $L$.
Notice that, to further account for the mentioned reward loss, whenever $l=0$ and adopt happens, we could consider $x$ in the reward function as 1-1 or $\frac{1}{2}$ in which case $y$ also should be increased by $\frac{1}{2}$. Although, we do not know whether the proof offering the most fee had $1$ unit reward or $\frac{1}{2}$ since this information is not included in the state space $(a,h,fork,p,l)$, this would only further increase the weight of reward split MDP in the final result. As we give the adversary the advantage of not losing any rewards, we ignore this situation. In any case, initiator proof heights are only $\frac{1}{L}$ of the total proof heights in the system.
D-A Simulations
To evaluate the performance of the reward split MDP and compare it with the Nakamoto consensus MDP of [5], we run the MDP models for $\gamma=0,0.2,0.4,0.5,0.6,0.8$ and $\alpha$ values ranges from $0.02$ to $0.48$ with $0.02$ spacing between them with $10^{-5}$ precision. It is clear that, reward splitting helps adversary when $\gamma$ is low as matching always gives a half reward when $(a,h,fork)=(1,1,relevant)$, whereas the adversarial relative reward in Nakamoto consensus increases with increasing $\gamma$ since matching succeeds more often. Thus, even when $\gamma=0.8$, the profitability threshold in our protocol stays around $\alpha\approx 0.28$.
@@ D-A Simulations
Appendix E Transaction Split Rule

### v2

@@ IV-B1 Proof Withholding
IV-B2 Numerical Evaluation
In Appendix F, we explain the MDP model with the appropriate state and action space as well as the objective function for our protocol. In Figure 5, we present the results of the MDP models together with the results of Section II where we analyzed the relative rewards of the existing parallel PoW protocols. We pick $\gamma=0.5$ for the MDP models and truncate the results of Section II before they rise above the honest behavior for the sake of neatness in the figures. For Nakamoto consensus, we run the MDP model of [5] and the modified model described in in Appendix F is run for our protocol with $10^{-5}$ precision where $\alpha$ values range from $0.02$ to $0.48$ with a $0.02$ spacing between them.
@@ IV-B2 Numerical Evaluation
In our protocol, matching always gives a half reward when $(a,h,fork)=(1,1,relevant)$ (see Appendix F) whereas the adversarial reward in Nakamoto consensus increases with increasing $\gamma$ since matching succeeds more often. Thus, in comparison to the Nakamoto consensus, our protocol is more resilient against selfish mining attacks when $\gamma\geq 0.3$ and resilience increases even more as $\gamma$ increases further, which we plot in Appendix F. On the other hand, when $\alpha>0.29$ as $\gamma$ drops below $0.3$, Nakamoto consensus becomes slightly more resilient. For the relative ratio in other values of $\gamma$, we refer the reader to Appendix F.
@@ IV-C Transaction Fee Split Rule
In Appendix G, we consider such an attacker that acts against the rule of picking the ledger offering the most fees to increase its expected transaction fee revenue. We show that, as long as $r$ is small enough, such an attack is usually not profitable. Moreover, even if profitable, the effect on the transaction fee rewards received by honest miners is minimal.
V Relation to existing works and limitations
@@ Appendix B Why Relative Reward Metric?
To understand why relative reward metric is the correct choice and why the normalized reward metric of [28, 29] is an illusion, consider the adversarial (honest resp.) miners who hold $\alpha$ ($1-\alpha$ resp.) fraction of the hashpower in the system. Let the total adversarial (honest resp.) reward per chain progress be denoted as $\mu$ ($\nu$ resp.). Clearly, an incentive compatible system that is not under an attack, should satisfy
@@
We will analyze this system in more detail. However, to give the gist of our argument in a few words, if such a system that is under attack is safe as claimed in [28, 29], then honest mining should be sustainable even if the adversary launches the attack. However, if honest mining is sustainable under the attack given the mining rewards and costs, then the adversarial mining is not only sustainable but also more profitable since adversary gets more relative rewards compared to its costs.
@@
$\frac{\nu_{attack}}{1-\alpha}<\frac{\nu_{attack}}{1-\rho}=\frac{\mu_{attack}}{\rho}<\frac{\mu_{attack}}{\alpha}.$ (10)
@@
$\frac{\mu_{attack}}{\alpha}>\frac{\mu_{attack}}{\rho}>\frac{\nu_{attack}}{1-\alpha},$ (12)
@@ Appendix C Proof Withholding Attack for Tailstorm and DAG-style Voting
Figure 7: Tailstorm withholding attack
@@ C-B Simulations
Appendix D Fork Choice Rule
The fork choice rule is to pick the chain that contains the most aggregate work at all times. Thus, if $L$ proofs are mined for a block and the adversary owns the valid ledger offering the most fee but shares the ledger with only some of the honest miners in the grace period, these honest miners try to create an initiator proof of the next block height referring to the adversarial ledger while other honest miners try to create an initiator proof referring to another ledger. However, this is not a threat for consensus mechanism since as soon as someone creates and shares the next initiator proof, every honest miner will pick that newly created proof and the ledger it refers to as the fork choice since it contains the most aggregate work. In such a situation, by not sharing the ledger with all honest miners in time, the adversary risks any potential extra income due to owning the ledger offering the most fee since the first initiator proof created next might refer to another ledger as the winning ledger.
On the other hand, if an honest miner owns the ledger offering the most fee and the adversary creates the first initiator proof for the next block height, honest miners will accept the adversarial initiator proof since it is the chain with most aggregate work. Notice, the adversarial initiator proof might refer to another valid ledger instead of the ledger offering the most fee. Such a situation does not create any issue from the perspective of the consensus since the the fork choice rule is always the same, i.e., pick the chain that contains the most aggregate work. From the perspective of the transaction fee revenues, in Section IV-C and Appendix G we show that such an attack is not profitable for the adversary.
Appendix E Analysis of the Double-spending Attack
E-1 The Attack
Essentially, we have to analyze the following. At block $h+1$, the miners start mining (voting) on ledgers that are created by miners of the proofs at block $h$. When an honest ledger of block $h$ receives $L$ votes at block $h+1$ and is part of the heaviest chain, the ledger is confirmed. See for example $\omega_{6}^{h}$ in Figure 3, which is confirmed when block $h+1$ is fully created. Any time after the honest ledger is confirmed, an adversarial ledger from block $h$, for example $\omega_{7}^{h}$ in Figure 3, is confirmed and replaces the honest ledger, if a heavier chain is published. Such a replacement is an instance of double-spending in our analysis.
Here, transaction fee fork choice rule does not matter as the most aggregate rule is always in place and we break ties in the favor of the adversary. $\Delta$ considered here can be seen as $2$ times the original $\Delta$ value in the protocol description since the adversary can avoid releasing its ledger if it offers the most fee resulting in doubling of a valid ledger propagation time. We ignore such a triviality here as it does not effect our analysis.
E-2 Analysis
Lemma 3
We skip proving this result as it essentially follows the same steps as bounded $\Delta$-delay model, by simply replacing blocks with proofs and $\Delta$ with $\delta$ in [30, Lemma 5]. We note that, since our protocol allows two parallel proofs to be combined, the balanced attacks are less likely to succeed and the rigged model assumption we make here is an extremely pessimistic one. Further, [7, 8, 9] show that more than two proofs being mined within a $\delta$ time is extremely unlikely, hence adversarial balanced attacks are less effective in our protocol.
$\lambda_{P}=L\lambda_{B},$ (19)
and
$p=1-q=\beta e^{-\lambda_{P}\delta}>\frac{1}{2}.$ (20)
Lemma 4
$S_{L}+S_{\delta_{L}}+S_{\Delta_{2}}.$ (21)
Lemma 5
The adversarial lead at $\tau$ can be upper bounded as
$M_{lead}\sim Geo\Big(1-\frac{q}{p}\Big).$ (22)
Proof:  We refer the reader to [30, Section 5.1] for the details of the proof. Essentially, we can assume $\tau$ to be large and the lead becomes the steady-state distribution birth-death process since it cannot go below zero, where the probability of a step forward is q=1pq=1-p, i.e., an adversarial arrival.  $\blacksquare$
Theorem 1
$P(M_{lead}+S_{L}+S_{\delta_{L}}+S_{\Delta_{2}}+M_{deficit}\geq L).$ (23)
Appendix F MDP Models for the Proposed Protocol
@@
To consider if this last case of expanded state space helps the adversary or not, we can first compare the relative rewards of MDP model of [5] and the reward split MDP we explained earlier. If the reward split MDP results in higher relative rewards for an $(\alpha,\gamma)$ pair, it makes no sense for the adversary to withhold the ledger with the most fees. On the other hand, if the reward split MDP results in less relative rewards, then it is in the interest of the adversary to launch this type of attack for initiator heights if it has a ledger offering the most fees.
We assume the adversary can always create a ledger with most fees offered, which is more than what the adversary could achieve. We implement this complicated situation where we have a state space of the form $(a,h,fork,p,l)$ with $l$ keeping track of the last common proof height. Even in the scenario where there is no reward loss, the result is approximately a weighted average between the reward ratio of the work in [5] (with approximate weight $1/L$) and reward split MDP (with approximate weight (L1)/L(L-1)/L). This is expected since the last case arises in approximately $\frac{1}{L}$ of the proof heights, i.e., the initiator heights. On the other hand, for $\gamma<0.3$, the result is the same as the reward ratio of reward split MDP since it is more profitable to have parallel proofs in this case. As the gain is minor even when the adversarial losses are ignored, we omit the simulation with expanded state space as the computational burden becomes very large with increasing $L$. In any case, initiator proof heights are only $\frac{1}{L}$ of the total proof heights in the system.
F-A Simulations
To evaluate the performance of the reward split MDP and compare it with the Nakamoto consensus MDP of [5], we run the MDP models for $\gamma=0,0.2,0.4,0.5,0.6,0.8$ and $\alpha$ values range from $0.02$ to $0.48$ with $0.02$ spacing between them with $10^{-5}$ precision. It is clear that, reward splitting helps adversary when $\gamma$ is low as matching always gives a half reward when $(a,h,fork)=(1,1,relevant)$, whereas the adversarial relative reward in Nakamoto consensus increases with increasing $\gamma$ since matching succeeds more often. Thus, even when $\gamma=0.8$, the profitability threshold in our protocol stays around $\alpha\approx 0.28$.
@@ F-A Simulations
Appendix G Transaction Split Rule